OpenStudy (watchmath):

Determine the convergence of the series \[\sum_{n=1}^\infty 1-\cos(1/n)\] (this is a problem from one of the student here)

OpenStudy (anonymous):

is this open for anyone to answer?

OpenStudy (watchmath):

sure

OpenStudy (anonymous):

i expanded cos(1/n) and rewrote the summand as: \[\sum_{n=1}^{\infty}(-1)^{n+1}/(2n)!n^{2n}\] which is an alternating series that is monotone decreasing, so it converges.

OpenStudy (watchmath):

you mean using the Taylor series? then shouldn't we have to sigma notation?

OpenStudy (watchmath):

two sigma notations

OpenStudy (anonymous):

oh yep i made a mistake.

OpenStudy (anonymous):

hey ty for posting it up the question but whats the answer is ∑n=1∞(−1)n+1/(2n)!n2n

OpenStudy (anonymous):

posted by rsvitale

OpenStudy (watchmath):

no, he made a mistake. I'll post the solution if nobody answer this :).

OpenStudy (toxicsugar22):

hi watchmath can u help me after him

OpenStudy (anonymous):

okay lol

OpenStudy (toxicsugar22):

watchmath can u help me

OpenStudy (anonymous):

by the way do you know how to find the general term of a series

OpenStudy (anonymous):

if given something like this 3,6,10,15

OpenStudy (anonymous):

i mean sequence

OpenStudy (watchmath):

you need to try to see the pattern and formulate it in general.

OpenStudy (anonymous):

so you basically guess

OpenStudy (watchmath):

No guessing is not the right word for it. It is observing some regularity :).

OpenStudy (watchmath):

Ok, let solve the series problem. By the double angle formula we know that 1- cos(2x)= 2sin^2(x) It follows that \(1-\cos(1/n)=2\sin^2(1/(2n))\) Recall that for positive x we have \(\sin x < x\) It follows that \(\sin^2(1/(2n))< (1/(2n) )^2\) But know the series \(\sum 1/(2n)^2\) is convergent (p-series with p=2). Hence \(\sum 2\sin^2(1/(2n))\) is convergent as well :).